OpenStudy (anonymous):

A right triangle has one side of length 7 and a hypotenuse of length 10. How long is the third side of the triangle?

Parth (parthkohli):

Use Pythagorean. \(\Large \color{Black}{\Rightarrow 7^2 + x^2 = 10^2 }\) \(\Large \color{Black}{\Rightarrow 49 + x^2 = 100 }\) \(\Large \color{Black}{\Rightarrow x^2 = 51 }\) \(\Large \color{Black}{\Rightarrow x = \sqrt{51} }\)